# Convergence and Generalization of Anti-Regularization for Parametric Models

### v1

@@ 2.7 Extension to Shallow MLPs (NTK Regime)
For 1–3 layer MLPs with sufficiently wide width and small learning rate, linearization around initialization holds, and the function space optimization under the NTK $K$ becomes:
@@
If $\lambda<\sigma_{\min}(K)$, then a minimizer exists, and Anti-regularization amplifies directions associated with small eigenvalues of $K$.
@@
The spectral argument for linear regression with $\hat{\Sigma}$ applies directly by replacing it with $K$. ∎
2.8 Stability Perspective and Necessity of $\lambda$ Decay
If the loss is $L$-smooth and the regularizer $R$ is $\alpha_{R}$-strongly convex, then the uniform stability of empirical risk minimization scales as
$\beta=O\!\left(\frac{L}{\,|S|\,(\mu-\lambda\,\alpha_{R})}\right),$
where $\mu$ denotes the strong convexity parameter of the empirical loss.
@@ 2.8 Stability Perspective and Necessity of $\lambda$ Decay
As $\lambda$ increases, $\mu-\lambda\alpha_{R}$ decreases, degrading the stability bound $\beta$. By adopting a decay schedule such as $|\lambda(n)|=|\lambda(n_{0})|\left(\tfrac{n_{0}}{n}\right)^{\alpha}$, we can ensure $\beta\to 0$. We recommend $\alpha\geq 1$ for regression and $\alpha\geq\tfrac{1}{2}$ for classification.
@@
Since the denominator is $|S|(\mu-\lambda\alpha_{R})$, decreasing $|\lambda|$ appropriately with increasing $n$ guarantees that the bound converges to zero. ∎
@@ 3.1 Overview
This study quantitatively evaluates the hypothesis that weak Anti-regularization (i.e., negative regularization) improves performance in small-sample (initial underfitting) regimes, and that gradually decaying the reward strength as more data accumulates leads to convergence toward zero regularization. To this end, we selected two distinct tasks—regression and classification—and systematically expanded dataset sizes. All comparisons were conducted under fixed architecture, preprocessing, and training budget, contrasting the baseline condition ($\lambda=0$) with Anti-regularization variants. To observe optimizer-specific effects, both Adam and SGD with momentum were employed. All evaluations were performed as paired comparisons across identical multi-seed setups to eliminate confounding from initialization and stochasticity.
@@ 3.2 Datasets and Tasks
To evaluate the behavior of Anti-regularization (AR) under varying signal-to-noise conditions and task complexities, we selected two regression and two classification datasets. For regression, we used the UCI Concrete Compressive Strength and UCI Airfoil Self-Noise datasets. Regression tasks directly reflect output scale and bias–variance trade-offs, making them well-suited to isolate the impact of AR on residual structure and sensitivity to output magnitude.
For classification, we employed MNIST and CIFAR-10. MNIST, with its high signal-to-noise ratio and simple structure, enables rapid decision boundary formation even with small samples, serving as a baseline to observe underfitting correction. In contrast, CIFAR-10 presents higher complexity and noise, with slower convergence and greater generalization capacity, making it a more challenging setting to assess AR’s effects in classification.
To avoid potential fairness concerns, we excluded datasets that contain sensitive attributes—such as gender or race—or their proxies.
@@ 3.3 Data Regimes and Sampling
As the central hypothesis concerns behavior across varying sample sizes, we defined five data regimes per task. For regression, we used $1\%$, $5\%$, $10\%$, $25\%$, and $50\%$ of the full dataset. For classification, we selected $0.1\%$, $0.2\%$, $0.3\%$, $0.5\%$, and $1\%$ relative to the full training set. Each subset was stratified based on the target distribution (regression) or class labels (classification).
All data regimes, optimizers, and regularization conditions were repeated with five fixed random seeds. We report the average performance and variability across these seeds.
@@ 3.4 Model and Optimization
To ensure fair comparison, we used identical architectures and training budgets across all conditions. Both Adam and SGD with momentum (m=0.9m{=}0.9) were employed to observe how Anti-regularization interacts differently with each optimizer.
Batch size, learning rate, and number of epochs were chosen based on standard practices per task and held constant between the Baseline and AR settings. To prevent excessive parameter updates, we imposed a trust-region constraint with radius $B_{\ell}$ on each linear layer. When the reward term induced updates that exceeded this boundary, we applied a projection operator and gradient clipping as safeguards.
@@ 3.5 Anti-regularization and Schedules
The initial intensity $\lambda_{0}$ of Anti-regularization was searched over a grid of $\{10^{-4},\,3{\cdot}10^{-4},\,10^{-3},\,3{\cdot}10^{-3},\,10^{-2}\}$. The effective intensity was then attenuated according to the number of training samples $N$ by the schedule:
@@
where $\alpha{=}1.0$ was used for regression tasks and $\alpha{=}0.5$ for classification. Hereafter, $\lambda$ denotes the current value after scheduling, while $\lambda_{0}$ refers to the initial value.
This design ensures that as more data become available, the effect of the reward term vanishes naturally, allowing the model to transition into a no-regularization baseline regime. All $\lambda_{0}$ candidates were trained under identical budgets, and the representative setting was chosen based on validation performance for final test evaluation. The baseline condition ($\lambda{=}0$) was trained and evaluated under the same protocol to enable direct comparison.
@@ 3.6 Evaluation Metrics and Diagnostics
For regression, we use RMSE, MAE, and $R^{2}$. For classification, accuracy and cross-entropy serve as primary metrics, and Expected Calibration Error (ECE) is included to assess practical reliability. To monitor the stability of the reward term and diagnose potential failure modes of the safety mechanism, we track the following internal indicators:
Output scale ratio $\rho$.
Let $Z$ be the matrix of raw predictions on the validation set at the final epoch—either regression outputs ($N\times 1$) or pre-softmax logits ($N\times K$) for classification. Define the scale as the Frobenius norm $\|Z\|_{F}$. Let $\rho$ denote the ratio of scales between Anti-regularization (AR) and Baseline under the same seed, split, and optimizer:
@@
We report $\rho$ at the final epoch only.
@@
This is the proportion of mini-batches in the final epoch where gradient clipping was actually applied. The clipping threshold $\tau$ is fixed across all runs.
@@
This is the proportion of mini-batches in the final epoch where the projection operator was invoked.
Triggering rule.
In regression, we define the safe operating zone as $r_{\mathrm{clip}}\in[0.05,\,0.40]$ and $\rho\in[0.9,\,1.3]$. In classification, the recommended bounds are $r_{\mathrm{clip}}\in[0.10,\,0.50]$ and $\rho\in[0.8,\,1.1]$. If either metric stays out of bounds for 2–3 consecutive epochs, we reduce $\lambda\leftarrow\lambda/2$ as a conservative fallback. Conversely, if $r_{\mathrm{clip}}<0.10$ and validation metrics improve, we apply a mild increase: $\lambda\leftarrow 1.5\,\lambda$. These rules are layered on top of the power-law decay schedule $|\lambda(n)|\propto n^{-\alpha}$.
@@ 3.7 Validation Protocol and Selection
For each combination of data fraction and optimizer, we train all $\lambda_{0}$ candidates under the same schedule and select the representative configuration based on the primary task metric—RMSE for regression and accuracy for classification—measured on the validation set. Only the selected configuration is officially reported on the test set. To mitigate selection bias, the Baseline ($\lambda=0$) is also retrained and evaluated using the same seeds and repetitions. All comparisons are aggregated via seed-wise pairing to ensure that differences are not confounded by initialization randomness.
@@ 3.8 Risk Monitoring and Stop Rules
To prevent failure modes in practical deployments where the reward term excessively amplifies model outputs, we continuously monitor $\rho$ and $r_{\mathrm{clip}}$ during training. If the final-epoch $\rho$ increases significantly relative to the Baseline, or if $r_{\mathrm{clip}}$ shows a sudden surge, the corresponding $\lambda_{0}$ is deemed unsuitable. Example thresholds include $\rho>1.3$ or $r_{\mathrm{clip}}>0.6$, in which case we either accelerate the decay exponent $\alpha$ or disable the reward term. For classification, a validation ECE increase greater than $0.02$ also triggers the same response. This rule serves as a practical instantiation of the theoretical safety condition; however, it was not applied in the reported experiments.
@@ 3.9 Statistical Treatment and Reporting
Given the small-data regimes where effects may be subtle, all conditions are evaluated with multiple random seeds. Primary comparisons report seed-wise paired means with standard errors, and, where necessary, supplemented by paired-sample statistical tests. Tables and main text summarize representative results, while the full $\lambda_{0}$ grid, per-seed raw metrics, $\lambda$ values, $\rho$, $r_{\mathrm{clip}}$, and $r_{\mathrm{proj}}$ logs are made publicly available in the appendix and repository to enable independent verification.
@@ 4.1.1 UCI Concrete Compressive Strength
Statistical comparisons in Table 2 confirm the directional consistency of AR’s effects, though 95% confidence levels were not reached under conservative testing with 32 seeds. Holm–Bonferroni adjusted $p_{t}^{\mathrm{Holm}}$ remained at $1.0$ across all blocks, and unadjusted $p_{t}$ and $p_{W}$ generally exceeded the significance threshold. However, the 95% confidence intervals narrowed quickly with increasing sample size (e.g., 50% RMSE CI: $[6.392,6.772]$), indicating that the practical gains observed at intermediate fractions are meaningful. This suggests that AR is particularly effective in mid-sized regimes, and statistical power may improve with increased replication or larger seed sets.
Stability and safety diagnostics in Table 3 support that the chosen AR configurations operated within safe zones. Gradient clipping was applied in 100% of updates across all fractions, ensuring numerical safeguards. The projection intervention rate ($r_{\mathrm{proj}}$) remained low to moderate (14–26%) for small to medium splits, but increased significantly at 25–50% (62.7, 74.9), reflecting more frequent activation of the trust-region constraint without harming convergence. The output scale diagnostic $\rho$ remained near zero at the $10^{-3}$ scale (only 4.3 at the most difficult 1% split), suggesting no signs of output explosion or divergence. Epoch counts and runtime scaled predictably with data size.
@@ 4.1.2 UCI Airfoil Self-Noise
Statistical inference in Table 5 shows that these directional improvements reach partial statistical support in the smallest-data regime. At 1%, the Wilcoxon test achieves significance for $R^{2}$ ($p_{W}=0.0413$), and approaches the threshold for RMSE ($p_{W}=0.0521$). In contrast, the paired $t$-test with Holm–Bonferroni correction did not reach significance at any fraction. For other splits, both tests recommend conservative interpretation. Nonetheless, confidence intervals narrowed with increasing sample size (e.g., 50% RMSE: $[3.301,3.644]$, $R^{2}$: $[0.707,0.764]$), indicating greater stability in measurement. In summary, AR’s effectiveness is most pronounced in the 1–5% range, and its statistical detectability may improve with more seeds or repetitions.
Stability and safety diagnostics in Table 6 support that the entire training process remained within safe bounds. The gradient clipping rate $r_{\mathrm{clip}}$ was 100.0% across all fractions, meaning all updates were protected by numerical safeguards. The projection intervention rate $r_{\mathrm{proj}}$ was relatively high at 1% (39.0), dropped to 6–16 in the 5–25% range, and rose slightly at 50% (9.5), suggesting that parameters stayed within safe regions as training data increased. The output scale diagnostic $\rho$ remained effectively zero (mostly 0.1, rising only to 1.2–1.8 at low data), indicating no signs of output explosion or divergence. Epoch counts and runtime scaled monotonically with data size, as expected.
@@ 4.2.1 MNIST
According to Table 7, the optimal configuration was $\lambda_{0}=0$ for most training data fractions, except for the 0.5% split, where the SGDM and $\lambda_{0}=10^{-3}$ combination showed a minor benefit (ECE 0.007-0.007, Accuracy +0.005+0.005). That is, Anti-regularization (AR) acted largely neutrally on this task, with only marginal gains in a narrow low-data regime. As the data fraction increased, both Accuracy and calibration (lower ECE) improved naturally, and even at the 1.0% split, the model converged to strong baseline performance without AR.
Statistical comparisons in Table 8 confirm this observation quantitatively. Even for the 0.5% split where AR was selected, $p$-values from both the $t$-test and Wilcoxon test exceeded conservative thresholds (i.e., not significant after Holm–Bonferroni correction). For other splits, AR was not selected ($\lambda_{0}=0$), making significance testing unnecessary (denoted by em dashes). Overall, the gains from AR on this dataset were minimal and not statistically substantiated.
Stability and safety diagnostics in Table 9 indicate that training proceeded stably within safe bounds across all splits. The projection intervention rate $r_{\mathrm{proj}}$ was 100.0 in every case, suggesting the trust-region constraint was active at every step. The gradient clipping rate $r_{\mathrm{clip}}$ increased with data size, ranging from 0–40%. The output scale diagnostic $\rho$ remained near zero for most splits; even at the 0.5% split where AR was active, $\rho$ was measured as 158.9 (at $10^{-3}$ scale) without showing divergence. Epochs and runtime followed a typical monotonic increase with dataset size.

### v2

@@ 3.7 Extension to Shallow MLPs (NTK Regime)
For 1–3 layer MLPs with sufficiently wide width and small learning rates, linearization in function space holds near initialization. This reduces to the NTK $K$ formulation:
@@
If $\lambda<\sigma_{\min}(K)$, then a minimizer exists, and Anti-regularization amplifies the directions corresponding to small eigenvalues of $K$.
@@
The spectral argument for linear regression with $\hat{\Sigma}$ applies directly when replaced by $K$. ∎
Remark 3 (Case when NTK is PSD).
Empirically, the NTK $K$ is often positive semidefinite (PSD), in which case $\sigma_{\min}(K)=0$. Then the safety region defined by $\lambda<\sigma_{\min}(K)$ is empty. To address this, one may ridge-augment the kernel as $K+\varepsilon I$ and take the limit as $\varepsilon\to 0$, or restrict to the positive eigenspace of $K$ and interpret $(K-\lambda I)^{-1}$ using the Moore–Penrose pseudoinverse. Under either treatment, the closed-form solution and stability arguments in the main text remain valid under the same assumptions.
3.8 Stability Perspective and the Necessity of Decreasing $\lambda$
When the loss is $L$-smooth and the regularizer $R$ is $\alpha_{R}$-strongly convex, the uniform stability of empirical risk minimization scales as
$\beta=O\!\Big{(}\frac{L}{\,|S|\,(\mu-\lambda\,\alpha_{R})}\Big{)},$
where $\mu$ denotes the strong convexity of the empirical loss.
@@ 3.8 Stability Perspective and the Necessity of Decreasing $\lambda$
As $\lambda$ increases, $\mu-\lambda\alpha_{R}$ decreases, and thus the bound on $\beta$ deteriorates. By adopting a schedule such as $|\lambda(n)|=|\lambda(n_{0})|\big{(}\tfrac{n_{0}}{n}\big{)}^{\alpha}$, one recovers $\beta\to 0$ as $n$ grows. For regression, $\alpha\geq 1$ is recommended, while for classification, $\alpha\geq\tfrac{1}{2}$ is sufficient.
@@
Since the denominator is $|S|(\mu-\lambda\alpha_{R})$, decreasing $|\lambda|$ sufficiently as $n$ increases ensures that the stability bound converges to zero. ∎
@@ 4.1 Overview
This study quantitatively evaluates the hypothesis that weak Anti-regularization (negative regularization) improves performance in small-sample regimes characterized by initial underfitting, while its reward strength decays toward zero as the sample size increases. To this end, we adopt a systematic design that spans both regression and classification tasks, with dataset sizes expanded step by step. All comparisons are conducted under fixed architectures, preprocessing pipelines, and training budgets, with Baseline ($\lambda=0$) and Anti-regularization conditions crossed against each other. To disentangle optimizer-specific effects, both Adam and SGD with momentum are employed. To ensure that observed differences do not arise from initialization or stochasticity, all settings are run with identical multiple seeds under paired comparisons.
@@ 4.2 Datasets and Tasks
To evaluate the behavior of Anti-regularization across varying signal-to-noise environments and task difficulties, we employed two regression and two classification datasets. For regression, we used UCI Concrete Compressive Strength and UCI Airfoil Self-Noise. Regression tasks directly reflect the impact of output scale and the bias–variance trade-off on performance, making them well-suited for isolating the effects of AR on residual structure and sensitivity to output scaling. For classification, we used MNIST and CIFAR-10. MNIST, with its high signal-to-noise ratio and simple structure, allows decision boundaries to form quickly even with few samples, serving as a baseline for assessing AR’s effect on correcting initial underfitting. CIFAR-10, in contrast, is more complex and noisy, converges more slowly, and offers greater room for generalization, thereby providing a more challenging setting for evaluating AR in classification. To avoid potential fairness concerns, datasets containing sensitive attributes such as gender or race, or their proxies, were excluded.
@@ 4.3 Data Regimes and Sampling
Since the core hypothesis concerns differences in behavior with respect to sample size, we considered $1\%$, $5\%$, $10\%$, $25\%$, and $50\%$ of the data for regression, and $0.1\%$, $0.2\%$, $0.3\%$, $0.5\%$, and $1\%$ of the training set for classification. Subsets were constructed using stratification by class or target distribution. All proportions, optimizers, and regularization conditions were repeated with five identical seeds, reporting both mean performance and variability.
@@ 4.4 Model and Optimization
For fair comparison, all conditions used identical architectures and training budgets. Optimizers included Adam and SGDM (m=0.9m{=}0.9), allowing us to disentangle optimizer-specific sensitivities to Anti-regularization. Batch size, learning rate, and number of epochs were set to conventional task-specific values and held fixed across both Baseline and AR conditions. To prevent excessive parameter updates, each linear layer was subject to a trust-region constraint of radius $B_{\ell}$. When updates from the reward term exceeded this boundary, a projection operator and gradient clipping were applied.
@@ 4.5 Anti-regularization and Schedules
The initial strength $\lambda_{0}$ of Anti-regularization is searched over the grid $\{10^{-4},\,3{\cdot}10^{-4},\,10^{-3},\,3{\cdot}10^{-3},\,10^{-2}\}$. The effective strength then decays with sample size $N$ as
@@
with $\alpha{=}1.0$ for regression and $\alpha{=}0.5$ for classification. Hereafter, $\lambda$ denotes the current coefficient under this schedule, while $\lambda_{0}$ refers to the initial value. This ensures that as data accumulates, the reward strength naturally decays toward zero, converging to the Baseline in the transition and saturation regimes. All $\lambda_{0}$ candidates are trained under identical budgets, and the representative configuration is selected based on validation performance for reporting test results. The Baseline ($\lambda{=}0$) is trained and evaluated under the same protocol for direct comparison.
@@ 4.6 Evaluation Metrics and Diagnostics
For regression, we use RMSE, MAE, and $R^{2}$. For classification, accuracy and cross-entropy serve as the primary metrics, with expected calibration error (ECE) additionally reported to assess practical reliability. To monitor the stability of the reward term and diagnose potential failure modes of safeguards, we record the following internal indicators.
Output-scale ratio $\rho$.
At the final epoch, we collect raw predictions (regression) or pre-softmax logits (classification) into a matrix $Z$ ($N\times K$ for classification, $N\times 1$ for regression). Scale is defined by the Frobenius norm $\|Z\|_{F}$. Comparing Anti-regularization with the Baseline under identical seeds, splits, and optimizers:
@@
We report $\rho$ only at the final epoch.
@@
The proportion of mini-batches in which gradient clipping was applied at the final epoch. The clipping threshold $\tau$ is fixed across all experiments.
@@
The proportion of mini-batches in which the projection operator was applied at the final epoch.
Operational trigger rules.
For regression, we recommend the ranges $r_{\mathrm{clip}}\in[0.05,\,0.40]$ and $\rho\in[0.9,\,1.3]$. For classification, we recommend $r_{\mathrm{clip}}\in[0.10,\,0.50]$ and $\rho\in[0.8,\,1.1]$. If either indicator leaves its recommended range for 2–3 consecutive epochs, $\lambda$ is conservatively halved. Conversely, if $r_{\mathrm{clip}}<0.10$ while validation metrics improve, $\lambda$ is increased modestly by a factor of 1.5. These rules are applied in conjunction with the power-law decay schedule $|\lambda(n)|\propto n^{-\alpha}$.
@@ 4.7 Validation Protocol and Selection
For each dataset fraction and optimizer combination, the $\lambda_{0}$ grid is trained under the same schedule, and a representative configuration is selected based on the primary task metric (RMSE for regression, accuracy for classification) on the validation set. Only the selected configurations are officially reported for test performance. To mitigate selection bias, the Baseline is retrained and evaluated with the same repetitions and seeds. All comparisons are aggregated using seed-paired means so that differences between conditions are not distorted by initialization effects.
@@ 4.8 Risk Monitoring and Stop Rules
To prevent failure modes in deployment where the reward term could cause excessive output amplification, we continuously monitor $\rho$ and $r_{\mathrm{clip}}$ during training. If the final-epoch $\rho$ increases excessively relative to the Baseline or if $r_{\mathrm{clip}}$ exhibits sharp growth, the corresponding $\lambda_{0}$ is deemed unsuitable. As illustrative thresholds, we adopt $\rho>1.3$ or $r_{\mathrm{clip}}>0.6$; in such cases, the decay exponent $\alpha$ is accelerated or the reward term is disabled. For classification, if the absolute increase in validation ECE exceeds $0.02$, the same adjustment is made. These rules operationalize the theoretical safety conditions into a practical protocol, though they were not applied to the reported experiments.
@@ 4.9 Statistical Treatment and Reporting
Given that effects may be small in low-data regimes, all conditions were repeated across multiple seeds. The main comparisons report seed-paired means and standard errors, with paired-sample tests provided as supplementary analysis when appropriate. Tables and main text summarize representative results, while the full $\lambda_{0}$ grid, per-seed raw scores, and logs for $\lambda$, $\rho$, $r_{\mathrm{clip}}$, and $r_{\mathrm{proj}}$ are released in the appendix and repository to enable independent verification.
@@ 5.1.1 UCI Concrete Compressive Strength
Statistical comparisons (Table 2) show that while the direction of effects was consistent, significance was not reached under the conservative 95% level once the variability across 32 seeds was accounted for. Within each split–optimizer block, Holm–Bonferroni adjusted $p\_t^{\mathrm{Holm}}$ remained 1.0, and unadjusted $p_{t}$ and $p_{W}$ generally exceeded the significance threshold. Nonetheless, the 95% confidence intervals narrowed rapidly with increasing sample size (e.g., RMSE interval $[6.392,6.772]$ at the 50% split), and the practically meaningful improvements observed in intermediate splits remain interpretively significant. This suggests that AR’s primary effective regime lies in the mid-sample range, with room for stronger statistical power under larger repetitions and seed counts.
Stability and safety diagnostics (Table 3) confirmed that the applied configurations did not exit the safe region. Gradient clipping rate ($r_{\mathrm{clip}}$) was 100% across all splits, indicating that every update was protected by numerical safeguards. Projection operator activation rate ($r_{\mathrm{proj}}$) remained modest (14–26%) at smaller and medium splits, but rose substantially (62.7, 74.9) at the 25–50% splits. This shows that trust-region constraints intervened more frequently as sample size grew, yet convergence remained stable without performance loss. The output-scale safety ratio $\rho$ was essentially near zero at the $10^{-3}$ scale (only reaching 4.3 at the most difficult 1% split), with no evidence of output explosion or divergence. The expected monotonic increases in epoch counts and runtime with larger sample fractions were also observed.
@@ 5.1.2 UCI Airfoil Self-Noise
Statistical comparisons (Table 5) show that these directional improvements have partial statistical support in the smallest-sample regime. At the 1% split, the Wilcoxon test reached the 5% significance level for $R^{2}$ ($p_{W}=0.0413$), and for RMSE it was borderline ($p_{W}=0.0521$). In contrast, paired $t$-tests under Holm–Bonferroni correction did not reach significance at the same split, and all other fractions required more conservative interpretation. Nevertheless, confidence intervals narrowed rapidly with increasing sample size (e.g., RMSE $[3.301,3.644]$, $R^{2}$ $[0.707,0.764]$ at the 50% split), confirming greater stability of measurement. In summary, AR’s primary effective regime lies in the low-data setting (1–5%), and increased repetitions and seed counts may enhance the statistical detectability of smaller improvements.
Stability and safety diagnostics (Table 6) confirm that training proceeded safely under the protection of safeguards. The gradient clipping rate $r_{\mathrm{clip}}$ was 100.0% across all splits, indicating that every update was constrained by numerical guardrails. Projection operator intervention rate $r_{\mathrm{proj}}$ was relatively high at the 1% split (39.0), dropped to lower levels (6–16) in the 5–25% range, and rose slightly again (9.5) at the 50% split. This suggests that parameters increasingly remained within the safe region as sample size grew. The output-scale ratio $\rho$ converged to nearly zero at the $10^{-3}$ scale (mostly 0.1, with values of 1.2–1.8 only in the smallest-sample regime), showing no signs of output explosion or divergence. Epoch counts and runtime scaled monotonically with data size in line with expectations.
@@ 5.2.1 MNIST
As shown in Table 7, the optimal setting was $\lambda_{0}=0$ for most training fractions, with the sole exception of the 0.5% split where the combination of SGDM and $\lambda_{0}=10^{-3}$ yielded a modest gain (ECE 0.007-0.007, Accuracy +0.005+0.005). In other words, Anti-regularization generally acted neutrally on this task, providing only minor improvements in the very narrow low-sample regime (0.5%). As the data fraction increased, both accuracy and calibration (lower ECE) improved naturally, and even at the 1.0% split the performance converged to high baseline levels without AR.
Statistical comparisons (Table 8) quantitatively confirm these observations. Even in the 0.5% split where AR was selected, the $t$-test and Wilcoxon test $p$-values exceeded conservative significance thresholds (not significant under Holm–Bonferroni correction). For all other splits, since the optimal configuration was already $\lambda_{0}=0$, statistical testing was not applicable (denoted by em dash). In summary, on this dataset AR’s direction of improvement was limited, and the observed small gains were not statistically substantiated.
Stability and safety diagnostics (Table 9) show that training proceeded stably under safety conditions across all splits. Projection operator intervention rate $r_{\mathrm{proj}}$ was reported as 100.0 in all fractions, indicating that trust-region constraints were activated at every step. Gradient clipping rate $r_{\mathrm{clip}}$ increased with data fraction, ranging from 0–40%. The output-scale safety ratio $\rho$ remained near zero in most cases, and although it reached 158.9 at the $10^{-3}$ scale in the 0.5% split where AR was active, no signs of divergence were observed. Epoch counts and runtime scaled monotonically with data size, consistent with expected patterns.